kittybasil:

\[\frac{3}{14}+\frac{10}{21}\div \left(\frac{3}{7}\right)\left(\frac{9}{4}\right)=\]

kittybasil:

What is the answer? Symbolab gives two: 58/81 and 9/17 The second one isn't in the original format though. ❓❓

Shadow:

This one looks real fun, sec

kittybasil:

Er I meant 19/7 for the second one. Typo.

kittybasil:

So they are as follows:

  • \(\frac{58}{81}\)
  • \(\frac{19}{7}\)

Shadow:

Wolfram gave 19/7

kittybasil:

Same but it has the incorrect format (compared to OG question)

Shadow:

What do you mean by original format?

kittybasil:

It's not written in the same format as the original given question.

kittybasil:

I did get the same answer while solving on my own but if the format doesn't match it's not the same problem/exercise is it?

Shadow:

I'm honestly confused, 19/7 is a fraction and this problem is about fractions.

Shadow:

Honestly what did you do at 3/14 + 40/81

logan23:

ho

Shadow:

ig I have to do the big #s

Shadow:

Okay I just solved it by hand and added parentheses around 3/7 and 9/4 in wolfram, for both I got 803/1134.

Shadow:

\[\frac{ 3 }{ 14 } + \frac{ 10 }{ 21 } \div ((\frac{ 3 }{7})( \frac{ 9 }{4 }))\]

Shadow:

\[\frac{ 3 }{ 7 } \times \frac{ 9 }{ 4 } = \frac{ 27 }{ 28 }\] \[\frac{ 10 }{ 21} \div \frac{ 27 }{ 28 } = \frac{ 10 }{ 21} \times \frac{ 28 }{ 27 } = \frac{ 10 }{ 3 } \times \frac{ 4 }{ 27 } = \frac{ 40 }{ 81 }\] \[\frac{ 3 }{14 } + \frac{ 40 }{ 81} = \frac{ 243 }{ 1134 } + \frac{ 560 }{ 1134 } = \frac{ 803 }{ 1134 }\]

kittybasil:

... wait what? Sorry I was multitasking @Shadow

kittybasil:

There's no parentheses around those two fractions, lol

Shadow:

Yeah I did that to show how I approached PEMDAS

Shadow:

I did the multiplication first, then division, then addition

Shadow:

Unless in your hw it has that fraction over the other fraction

kittybasil:

The answer in the book is \[\frac{19}{7}\]

kittybasil:

Which is what I got but when I put it into Symbolab it comes up as \[\frac{58}{81}\]

Shadow:

K then I guess it has to be done division first.

kittybasil:

I legit have no idea where you got that from tho lol

kittybasil:

Here's the one that has the "proper format" --

1 attachment
Shadow:

Yeah did the division first and got 19/7

Shadow:

I guess the division puts emphasis since that fraction is under the second one, then you multiply it with the 4th, then add the product with the first fraction.

kittybasil:

Under? ðŸĪ”

Shadow:

\[\frac{ \frac{ 10 }{ 21} }{ \frac{ 3 }{ 7 } }\]

kittybasil:

Oh so you mean it ends up looking like this?\[\frac{3}{14}+\frac{\frac{10}{21}}{(\frac{3}{7})(\frac{9}{4})}\]

kittybasil:

-- confusion

Shadow:

I'll just show my steps

kittybasil:

Is what I put wrong?

kittybasil:

The last equation thing I eman

kittybasil:

mean* Argh typos

Shadow:

\[\frac{ \frac{ 10 }{ 21} }{ \frac{ 3 }{ 7 } }\ = \frac{ 10 }{ 21} \times \frac{ 7 }{ 3 } = \frac{ 10 }{ 3 } \times \frac{ 1 }{ 3 } = \frac{ 10 }{ 9 }\] \[\frac{ 10 }{ 9 } \times \frac{ 9 }{ 4 } = \frac{ 10 }{ 1 } \times \frac{ 1 }{ 4 }= \frac{ 10 }{4 }\] \[\frac{ 3 }{ 14 } + \frac{ 10 }{ 4 }= \frac{ 6 }{ 28 }+ \frac{ 70 }{ 28 } = \frac{ 76 }{ 28 } = \frac{ 19 }{ 7 }\]

Shadow:

This way gives the correct answer.

Shadow:

So you'd put parentheses around the middle fractions. Apparently if you do the last two fractions first then divide, it gives the incorrect answer???

kittybasil:

Apparently it has to do with PEMDAS being left-right if the P and E are nonexistent. Not sure -- does the parenthetical part at the end not count as P?

Shadow:

I think I see the problem with Symbo, it's putting 3/14 on the top when it should be to the side

Shadow:



kittybasil:

\[\frac{3}{14}+\frac{10}{21}\div \left(\frac{3}{7}\right)\left(\frac{9}{4}\right)=\]\[\frac{3}{14}+\frac{10}{21}\times(\frac{7}{3})(\frac{4}{9})\]This would work too, right?

Shadow:

Yes, that's what I did but in that instance you'd have to do the middle fractions after flipping the reciprocal, since you're doing the division, then you can go to the last fraction.. then add.

kittybasil:

Testing.\[\frac{3}{14}+\frac{10\times7\times4}{21\times3\times9}\]\[\frac{3}{14}+\frac{10\times4}{3\times3\times9}\]

kittybasil:

Sorry, too lazy to put equals signs because I'm afraid my train of thought will derail. 😂\[\frac{3}{14}+\frac{40}{9\times9}\]\[\frac{3}{14}+\frac{40}{81}\]...Okay maybe it wouldn't work.

Shadow:

Yeah then you're doing what I did earlier, with 1134 in the denominator.

kittybasil:

\[\frac{3}{14}+\frac{10}{21}\div \left(\frac{3}{7}\right)\left(\frac{9}{4}\right)=\]\[\frac{3}{14}+\frac{10}{21}\times(\frac{7}{3})(\frac{9}{4})\]Then this?

kittybasil:

Basically removing the \(\frac{9}{4}\) as it's not part of the current operation being used

kittybasil:

And then integrating it back in later. I think the issue is that the method getting the giant fraction result is because the last 2 fractions are still essentially getting multiplied first Welp

Shadow:

Yeah I think book is assuming parentheses around 10/21 div 3/7

Shadow:

Because you need to do that first to return 19/7

Shadow:

When we multiply across we get a nasty fraction.

kittybasil:

There was actually a consecutive question that asked you to explain the order of operations you'd undertake. It didn't say that, but it did say PEMDAS without P & E = left-to-right.

kittybasil:

It didn't say "parentheses" in the answer key* oops

Shadow:

Honestly I thought it didn't make a difference when there's multiplication and division. Thought they had the same priority. I don't remember left to right lol.

Shadow:

But that would explain the priority.

kittybasil:

I wonder why Symbolab put that answer for a question written in the original format, i.e. \[\frac{3}{14}+\frac{10}{21}\div \left(\frac{3}{7}\right)\left(\frac{9}{4}\right)=\]

kittybasil:

But the correct answer had a different way of writing entirely

kittybasil:

I mean:

1 attachment
kittybasil:

Was it to emphasize the special rule of PEMDAS in this scenario...? Confusin

Shadow:

Yes it's taking it left to right.

Shadow:

So when they have the same priority like M/D or A/S, you go left to right

kittybasil:

Yeah, but I don't see why they couldn't have made that format step 2. I legitimately thought that was something else entirely

Shadow:

Can't just do the multiplication first.

kittybasil:

It's the parentheses. Trick shot. ðŸĨī

kittybasil:

I mean I probably did the thing with division-first but I had a full-on argument with someone over the fallibility of SL oops. 😂😂

Shadow:

Well technically there's no parentheses, it's just saying according to PEMDAS + left to right, need to do the division first, so it's pseudo parentheses if you mean as a term for priority.

Shadow:

Oh

Shadow:

Lmao

kittybasil:

Yeah, pseudo. Argh, this was so stupid xD

Shadow:

Yeah well honestly I learned something here so ig I'm grateful you asked.

Shadow:

I haven't done an actual math class in like 2yrs

kittybasil:

Feels more like a hardcore review LOL

kittybasil:

Because technically those parentheticals mean multiplication yet look like actual Ps from PEMDAS. ðŸĪĶðŸŧ‍♀ïļ

kittybasil:

Guess Symbolab got confounded ðŸĨī

kittybasil:

I mean, I get it, the issue was just me thinking that Symbolab had to be right so I was trying to prove it Whoops

Shadow:

I did some quick math to illustrate the rule: \[\frac{ a }{ b } \div \frac{ c }{ d } \times \frac{ e }{ f } = \frac{ a }{ b } \times \frac{ d }{ c } \times \frac{ e }{ f } = \frac{ ade }{ bcf }\] \[\frac{ a }{ b } \div \frac{ c }{ d } \times \frac{ e }{f } = \frac{ a }{ b } \times \frac{ df }{ ce } = \frac{ adf }{ bce }\] First one I did division of the first two terms first. Second one I did multiplication of the last two terms first. They return different results.

Shadow:

And left to right aka div first in this case is the right way.

Shadow:

Maybe Symbolab isn't applying the left to right rule? idk

kittybasil:

Honestly it's a literal WTH moment like. If it had gotten the 1134-denominator fraction I would have accepted it as a glitch in the system, but not THIS.

kittybasil:

I have no idea where this came from. It was so outrageous I started questioning if I had done something wrong lol

Shadow:

ig, but sometimes your teacher gives you problems with unpretty solutions. I don't know your prof so I just accepted the result, but I've learned now I gotta do left to right lul.

Shadow:

I think we got it though.

kittybasil:

Oh it's not me, it was someone else

Shadow:

The math engines are just computing things with an odd priority

kittybasil:

I'm way past this level but Symbolab threw me for a loop. I need to stop believing in the infallibility of technology lol.

Shadow:

I think the problem is the divisor, they assume 3/14 is on top with 10/21, and that 3/7 and 9/4 are on the bottom. Or some other other combination where 3/14 is outside.

Shadow:

See what I mean?

kittybasil:

Rather weird bc there aren't ANY parenthetical indicators. Like... what even.

kittybasil:

Yeah so by hand I would've had this, but I thought I got auto-corrected WHOOPS.

Shadow:

Yeah ig the engines are sometimes weird.

kittybasil:

Entire point of this thread: don't trust the internet LOL Alright thanks for putting up with that xD

Shadow:

This is why I usually put parentheses around my fractions, so they don't spill into each other.

kittybasil:

Too much to ask for a medal? ðŸĨī

kittybasil:

🙏ðŸŧ

kittybasil:

I feel like I ran into a similar issue of some other algebraic principle last week. I should go look for it and figure out why there was a disparity in that one...

kittybasil:

Found it

kittybasil:

Oh wait, that's a word problem. ðŸĪŠ

kittybasil:

Anyway thanks for helping me figure out how faulty modern internet can get. I'll stop pestering you with pings now

Shadow:

It was a pleasure.